OpenStudy (anonymous):

Which equation represents the line that passes through (–6, 7) and (–3, 6)? y = –x + 9 y = –x + 5 y = –3x – 11y y = –3x + 25

OpenStudy (anonymous):

Get the gradient of the line first, then put it into y=mx+b form.

OpenStudy (anonymous):

\[Gradient, m = (\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} })\]

OpenStudy (anonymous):

so -6 - -3 and 7-6?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

so -9,1

OpenStudy (anonymous):

or positive 9

OpenStudy (anonymous):

Lemme show you

OpenStudy (anonymous):

\[\frac{ 6-7 }{ -3-6 } = \frac{ 1 }{ 9 }\]

OpenStudy (anonymous):

So you put it into the equation as \[y = \frac{ 1 }{ 9 } x +c\]

OpenStudy (anonymous):

now you plug in a point like (-6,7) into the equation. \[7 = \frac{ 1 }{ 9 } (-6) +c\] And solve for c

OpenStudy (anonymous):

oh ok i see! Thank you!